Points A, B, C, D are collinear and positioned in that order. Find the length indicated.
AD=6x-385 AC=9x-735 BD=99 BC=10x-902. Find AC.

Answer by greenestamps(13200)



(1) AB = AD-BD = (6x-385)-99
(2) AB = AC-BC = (9x-735)-(10x-902)

(6x-385)-99 = (9x-735)-(10x-902)

(3) solve for x
(4) AC = 9x-735